# Reconstruction Alignment Improves Unified Multimodal Models

### v2

@@ Appendix E Limitations and Future Work
Architecture-specific constraints. RecA exhibits limited effectiveness across certain architecture: (I) Show-o with image tokenization. For models employing discrete image tokenizers as vision encoders for understanding tasks, we observe a tendency toward learning trivial one-to-one mappings. During training, Show-o demonstrates abrupt convergence behavior with cross-entropy loss dropping to near-zero values at specific training timesteps, suggesting potential overfitting phenomena. Future research could address this through regularization strategies such as input perturbation (e.g., image blurring, noise injection). (II) BLIP-3o. As illustrated in Figure 10, BLIP-3o possesses inherent image reconstruction capabilities. Application of RecA fails to improve reconstruction capability and text-to-image generation capability. We hypothesize that BLIP-3o’s pre-training already incorporates reconstruction objectives, rendering additional reconstruction training counterproductive.
@@
Appendix F Generated Captions
F.1 Captions for Figure 2
@@ F.2 Captions for Figure 8
Appendix G More Qualitative Results on Image Editing
We provide additional qualitative comparisons in this section. Figure 13 presents a comparison between the RecA post-trained BAGEL and previous SOTA models, ICEdit (Zhang et al., 2025c), FLUX-Kontext (Labs et al., 2025), and GPT-4o-Image. Compared with ICEdit and FLUX-Kontext, our method demonstrates stronger instruction-following capability. In contrast to GPT-4o-Image, our method exhibits superior identity preservation and background fidelity. Figure 12 and 11 showcase consistent improvements in semantic consistency, instruction following, and visual quality preservation across various editing tasks including background modification, style transfer, and object manipulation.
@@ Appendix G More Qualitative Results on Image Editing
Appendix H More Qualitative Results on Text-to-Image Generation

### v3

@@ Appendix F Limitations and Future Work
Architecture-specific constraints. RecA still exhibits limited effectiveness on BLIP-3o. As illustrated in Figure 14, BLIP-3o possesses inherent image reconstruction capabilities. Application of RecA fails to improve reconstruction or text-to-image generation quality, likely because BLIP-3o’s pre-training already incorporates strong reconstruction objectives, rendering additional reconstruction training counterproductive.
@@
Appendix G Generated Captions
G.1 Captions for Figure 2
@@ G.2 Captions for Figure 8
Appendix H More Qualitative Results on Image Editing
We provide additional qualitative comparisons in this section. Figure 17 presents a comparison between the RecA post-trained BAGEL and previous SOTA models, ICEdit (Zhang et al., 2025c), FLUX-Kontext (Labs et al., 2025), and GPT-4o-Image. Compared with ICEdit and FLUX-Kontext, our method demonstrates stronger instruction-following capability. In contrast to GPT-4o-Image, our method exhibits superior identity preservation and background fidelity. Figure 15 and 16 showcase consistent improvements in semantic consistency, instruction following, and visual quality preservation across various editing tasks including background modification, style transfer, and object manipulation.
@@ Appendix H More Qualitative Results on Image Editing
Appendix I More Qualitative Results on Text-to-Image Generation
@@
Appendix J The Use of Large Language Models (LLMs)
During the preparation of this paper, we used large language model (LLM) only as a writing assistant to check grammar, spelling, and to polish the clarity of expression. The LLM was not involved in the research design, experiments, or analysis, and the authors take full responsibility for all content of this work.